# Contextualizing Recommendation Explanations with LLMs: A User Study

### v1

@@ Abstract
Large language models (LLMs) are increasingly prevalent in recommender systems, where LLMs can be used to generate personalized recommendations. Here, we examine how different LLM-generated explanations for movie recommendations affect users’ perceptions of cognitive, affective, and utilitarian needs and consumption intentions. In a pre-registered, between-subject online experiment (N=759) and follow-up interviews (N=30), we compare (a) LLM-generated generic explanations, and (b) LLM-generated contextualized explanations. Our findings show that contextualized explanations (i.e., explanations that incorporate users’ past behaviors) effectively meet users’ cognitive needs while increasing users’ intentions to watch recommended movies. However, adding explanations offers limited benefits in meeting users’ utilitarian and affective needs, raising concerns about the proper design and implications of LLM-generated explanations. Qualitative insights from interviews reveal that referencing users’ past preferences enhances trust and understanding but can feel excessive if overused. Furthermore, users with more active and positive engagement with the recommender system and movie-watching get substantial gains from contextualized explanations. Overall, our research clarifies how LLM-generated recommendations influence users’ motivations and behaviors, providing valuable insights for the future development of user-centric recommender systems, a key element in social media platforms and online ecosystems.
@@ 1 Introduction
The emergence of large language models (LLMs) presents potential for improving these recommendation explanations (Wu et al. 2024; Hou, Leach, and Huang 2024). Some studies have adopted LLMs to produce more natural and human-like explanations. However, current efforts often rely on generic explanation templates (Li, Zhang, and Chen 2020). These generic explanations overlook users’ contexts or preferences, thereby limiting the benefits of the explanations and underutilizing the LLMs’ reasoning capabilities. Existing research seldom provides a comprehensive understanding of how users perceive the received recommendations. Current work also rarely examines whether positive user perceptions translate into users’ consumption intentions.
Motivated by this gap, our study adopts established explanation measurement standards (Balog and Radlinski 2020) and interprets user responses through the lens of the Uses and Gratifications (U&G) framework (Katz, Blumler, and Gurevitch 1973; Lin et al. 2013). The U&G framework categorizes users’ diverse needs, including the desire for understanding and knowledge acquisition (Cognitive needs), the need for emotional connection and engagement (Affective needs), and the focus on practicality and task-oriented goals (Utilitarian needs). This framework provides a valuable perspective to understand how users actively select, consume, and engage with the received information. Our goal is to examine how different LLM-generated recommendation explanation types resonate with specific user needs and, in turn, influence their consumption intentions. We design the following research question:
How do different LLM-generated explanations respond to users’ diverse motivational needs (cognitive, affective, utilitarian) and affect their consumption intentions?
@@
In this study, we conducted a pre-registered, between-subject online experiment involving $N=759$ subjects from the U.S. in the context of movie recommendations. [^1] The experimental protocol was pre-registered at OSF: https://osf.io/j46vx/?view_only=dc2b1f641b33407ab8168736a783055a By drawing upon the U&G framework, we analyze users’ perceptions and intentions towards three types of recommendations generated by GPT-4 Turbo: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE). We also drew upon user profile information and conducted follow-up interviews ($N=30$) to interpret our results more deeply. To ensure the robustness of our study design and the generalizability of the results, we conducted supplementary experiments by using a different LLM (Claude 3.5 Sonnet) with an additional 140 users from the U.S. Figure 1 provides an overview of our study flow.
@@ 2 Related Work
2.1 Explanations in Recommender Systems
@@
In social media and online communities, recommendations help users to find items aligned with their preferences. Initial research centered on improving accuracy (Ma et al. 2023; Wu et al. 2023), but recent work emphasizes explainability to foster trust, transparency, and user satisfaction (Burkart and Huber 2021). Early approaches integrated explainability into model architectures, such as explainable matrix factorization leveraging user similarities (Abdollahi and Nasraoui 2017), tree-enhanced embeddings (Wang et al. 2018), and attention-based neural models (Seo et al. 2017; Chen et al. 2018). More recent efforts have expanded explanation strategies, ranging from visual (Tsai and Brusilovsky 2019) and comparative methods (Le and Lauw 2021) to sentiment-oriented narratives (Chen, Yan, and Wang 2019), as well as traditional item-, content-, and popularity-based explanations (Sarwar et al. 2001; Pazzani and Billsus 2007; Yang 2016). These approaches highlight the importance of making recommendations more understandable, interactive, and socially aligned to support user trust and transparency across online communities.
Beyond these approaches, LLMs offer new opportunities for generating better explanations within the dynamic, user-driven landscapes (Mohanty 2024). By producing natural language explanations, LLMs can improve user satisfaction and trust, especially in domains where social context and user interaction are central (Chang, Harper, and Terveen 2016). For instance, LLMs have been used to create more detailed item descriptions (Acharya, Singh, and Onoe 2023) or to enable ChatGPT-based recommenders to generate user-friendly explanations (Deldjoo 2024). Recent research suggests that more contextualized, user-centered explanations could enhance trust and engagement (Silva et al. 2024; Lyu et al. 2024), reinforcing the value of LLM-driven strategies that align closely with the evolving social and informational needs of online communities.
2.2 User Perceptions of Recommendation Explanations
Users increasingly rely on recommender systems to navigate overwhelming volumes of information. Research grounded in the U&G framework highlights that individuals engage with media to fulfill specific needs, which can be categorized as affective, cognitive, and utilitarian motivations (Katz, Blumler, and Gurevitch 1973; Korhan and Ersoy 2016). These motivations directly influence consumption intentions, as users are more likely to engage with media or systems that effectively address their various needs. High-quality explanations within recommender systems can significantly influence user engagement, satisfaction, and decision-making, particularly in socially dynamic environments where these systems satisfy informational and social needs (Mandl et al. 2011; Jameson et al. 2015). For instance, explanations can enhance perceived usefulness and trust, aligning with motivation needs such as ease of use and task efficiency (Korhan and Ersoy 2016). Extant research also offers valuable insights into factors shaping user perceptions. For example, evidence shows that users may prefer human-generated recommendations, particularly for hedonic products (Wien and Peluso 2021). Similarly, explanations reflecting user feedback and personalization resonate with individuals who seek tailored, value-reinforcing content (Lu et al. 2023). Differences in user profiles further emphasize the importance of detailed and specific explanations, particularly for users with high cognitive engagement who value transparency and trust (Chatti et al. 2022).
Natural language explanations further enable systems to fulfill diverse needs, such as reducing cognitive effort, fostering trust, and enhancing community integration (Chang, Harper, and Terveen 2016). In parallel, researchers have developed evaluation metrics to address these user-centric demands beyond traditional measures like BLEU and ROUGE (Papineni et al. 2002; Lin 2004). User-centered metrics now consider dimensions such as effectiveness, persuasiveness, satisfaction, and trust, emphasizing user needs in complex, community-driven contexts (Zhou et al. 2021; Ehsan et al. 2019; Hoffman et al. 2023).
@@ 2.3 Research Gap
While the importance of recommendation explanations is widely acknowledged, research seldom examines how users interpret and respond to them. Furthermore, extant studies often rely on generic templates and have not fully explored the potential of LLM-generated contextualized explanations to address users’ cognitive, affective, and utilitarian needs (see Table 1).
To address this gap, our study provides the first comprehensive user evaluation of LLM-generated contextualized explanations through the lens of the U&G framework. By examining how these explanations resonate with users’ multifaceted needs and influence their intentions, we aim to offer deeper insights into designing recommendation explanations that are more aligned with user motivations.
@@ 3.1 Overview
We conducted an online between-subject experiment, where subjects ($N=759$) were randomly assigned to one of three experimental conditions: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE).
@@ 3.2 Subjects
Initially, 1401 subjects completed the survey. We discarded responses from subjects failing an initial attention check. Subsequently, following prior research (Pennycook, Cannon, and Rand 2018) and our pre-registration, a subset of subjects were excluded for three reasons: (i) subjects who self-reported to have answered randomly, (ii) failed the second attention check, (iii) responded excessively fast (i.e., the duration was less than around 6 seconds for responding to the user perceptions). These users still received remuneration after the exclusion. Eventually, due to our strict exclusion criteria, 759 of the 1,401 subjects qualified, exceeding our estimated sample size of 750 to ensure the results achieve the desired statistical power.
@@ 3.3 Materials
We selected GPT-4 Turbo to generate movie recommendations for the main experiment. We conducted an additional experiment with Claude 3.5 Sonnet to validate the robustness of our study design and the generalizability of our findings (see supplementary materials). We precompiled all recommendations and explanations in a corpus to ensure that there was no variability in the outcomes across different LLM calls.
To create the corpus, we first selected 10 movies with diverse topics from the IMDb Top 250 list [^4] https://www.imdb.com/chart/top/ to compile a movie shortlist $\mathcal{M}$. These movies were chosen based on their high ratings or popularity. We ensured that the shortlist had broad coverage of user viewing preferences. Subsequently, we implemented a sequential recommendation strategy following the algorithm in Table 5 in Appendix A. With this strategy, sujects with the same viewing history and assigned conditions received the identical recommendation..
@@
We followed previous practices (Giray 2023) when designing the prompts. Table 3 shows examples of contextualized and generic explanations. We can see a clear difference: while both explanations incorporate various elements (e.g., feature, popularity, similarity) to support the recommendations, the contextualized explanations naturally integrate the recommended movies with the user’s viewing history, explicitly highlighting how the user’s inputs shape the recommendation. Meanwhile, the generic explanations focus exclusively on the recommended movies. All generated movies were cross-checked with the IMDB dataset [^5] https://developer.imdb.com/non-commercial-datasets/ to ensure no hallucinations from the LLMs. Additionally, the corpus was validated by two experts with strong expertise in the field to ensure that all explanations used in the experiment were consistent and of fair quality.
@@ 3.4 Procedure
Afterward, we randomly reached out to $N=30$ users for the follow-up interviews. [^6] Detailed interview questions in Appendix B. We integrated their responses to help interpret the quantitative analysis results and investigate how explanations affect user motivational needs.
@@ 3.5 Analyses
We conducted analyses based on user perceptions in the survey. Each participant could provide up to 10 ratings, resulting in a total of 2,732 ratings from 759 participants in the main experiment (916 in NE, 898 in GE, and 918 in CE). We used hypothesis testing to identify differences in perceptions and consumption intentions across recommendation types, followed by mixed-linear regression to examine how user characteristics explain variation in both variables. Additionally, we conducted follow-up interviews to further interpret the quantitative results.
Hypothesis testing: We hypothesized that there are significant differences in the perceptions and consumption intentions between users receiving different recommendations. We conducted hypothesis testing on the four dependent variables of interest: Cognitive, Affective, Utilitarian, and UserIntent across the different conditions. Since the data did not meet the normality assumption (Lilliefors 1967), we used a Mann–Whitney test with Holm correction to identify pairwise differences between conditions (McHugh 2011). Later, we use $\mu$ to report the mean.
@@
$\emph{Y}_{ij}$ $=\beta_{0}+\beta_{1}\emph{Treatment}_{j}+\beta_{2}\emph{Age}_{j}+\beta_{3}\emph{Gender}_{j}+\beta_{4}\emph{Education}_{j}$
$\quad+\beta_{5}\emph{MovieConsumption}_{j}+\beta_{6}\emph{RecSysFrequency}_{j}$ (1)
$\quad+\beta_{7}\emph{RecSysUtility}_{j}+\beta_{8}\emph{DisplayOrder}_{j}+u_{0j}+\epsilon_{ij}$
where ${Y}_{ij}$ denotes the dependent variables (i.e., Cognitive, Affective, Utilitarian, and UserIntent) for the $i$-th observation of the $j$-th subject, with intercepts $\beta_{0}$, subject-level random effects $u_{0j}$, coefficients $\beta_{1}$ through $\beta_{8}$, and an error term $\epsilon_{ij}$.

### v2

@@ Abstract
Large language models (LLMs) are increasingly integrated into recommender systems to generate natural language explanations for contextualized content. This study examines how different types of LLM-generated explanations shape user perceptions and behaviors in the context of movie recommendations. In a pre-registered, between-subject online experiment (N=759) and follow-up interviews (N=30), we compare LLM-generated (a) generic explanations, and (b) contextualized explanations. We find that contextualized explanations better address users’ cognitive and affective needs, and increase their intention to watch the recommended movies. Insights from interviews reveal that contextualization fosters trust but depends heavily on alignment with users’ genre and narrative preferences. Notably, including more historical movie references improves user perceptions and intent, though overly detailed explanations can feel redundant. Users with higher engagement or trust in the system benefit most from contextualization. These findings highlight the potential of LLMs to enhance user-centered recommendation experiences. We offer practical guidance for designing explanation strategies that balance personalization and cognitive effort to improve relevance, trust, and engagement across social media and entertainment platforms.
@@ 1 Introduction
The emergence of large language models (LLMs) presents potential for improving recommendation explanations (Wu et al. 2024; Hou, Leach, and Huang 2024). Some studies have adopted LLMs to produce natural and human-like explanations. However, current efforts often rely on generic explanation templates (Li, Zhang, and Chen 2020). These generic explanations overlook users’ contexts or preferences, thereby limiting the benefits of the explanations and underutilizing the LLMs’ reasoning capabilities. Existing research seldom provides a comprehensive understanding of how users perceive the received recommendations. Current work also rarely examines whether positive user perceptions translate into users’ consumption intentions.
Motivated by this gap, our study adopts established explanation measurement standards (Balog and Radlinski 2020) and interprets user responses through the lens of the Uses and Gratifications (U&G) framework (Katz, Blumler, and Gurevitch 1973; Lin et al. 2013). The U&G framework categorizes users’ diverse needs, including the desire for understanding and knowledge acquisition (Cognitive needs), the need for emotional connection and engagement (Affective needs), and the focus on practicality and task-oriented goals (Utilitarian needs). This framework provides a valuable perspective to understand how users actively select, consume, and engage with the received information. Our goal is to examine how different LLM-generated recommendation explanation types resonate with specific user needs and, in turn, influence their consumption intentions. Therefore, we designed the following research question:
How do LLM-generated explanations respond to users’ diverse motivational needs (cognitive, affective, utilitarian) and affect their consumption intentions?
@@
In this study, we conducted a pre-registered, between-subject online experiment involving NN=759 subjects from the U.S. in the context of movie recommendations. [^1] The experimental protocol was pre-registered at OSF: https://osf.io/j46vx/?view_only=dc2b1f641b33407ab8168736a783055a By drawing upon the U&G framework, we analyze users’ perceptions and intentions towards three types of recommendations generated by GPT-4 Turbo: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE). [^2] We focus on real-world recommendation tasks based on viewing histories without relying on explicit or private user preferences (e.g., as in cold-start problems of recommendation websites). As such, we use the term contextualized rather than personalized throughout the paper. This terminology aligns with related work (Bar, Maarouf, and Feuerriegel 2024) and reflects practical usage on platforms such as IMDb, Yelp, and Google Maps, where recommendations are often informed by user activity (e.g., viewing or browsing history) but not by direct feedback like ratings. We also drew upon user profile information and conducted follow-up interviews (NN=30) to interpret our results more deeply. To ensure the robustness of our study design and the generalizability of the results, we conducted supplementary experiments by using a different LLM (Claude 3.5 Sonnet) with an additional 140 users from the U.S. Figure 1 provides an overview of our study flow.
@@ 2 Related Work
2.1 LLM-Generated Explanations in RecSys
@@
In social media and online communities, recommendations help users find items aligned with their preferences. Initial research centered on improving predictive accuracy (Ma et al. 2023; Wu et al. 2023). However, growing attention to user trust and transparency has driven a shift toward explainable recommendations (Burkart and Huber 2021). Initial research centered on improving accuracy (Ma et al. 2023; Wu et al. 2023), but recent work emphasizes explainability to foster trust, transparency, and user satisfaction (Burkart and Huber 2021). Traditional recommendation techniques (e.g., matrix factorization with interpretable layers (Abdollahi and Nasraoui 2017) or attention-based neural models (Seo et al. 2017; Chen et al. 2018)) integrated explainability into model architectures.
The advent of LLMs has made it easy to generate explanations in natural language more flexibly. Building upon template-based methods (Sarwar et al. 2001; Pazzani and Billsus 2007; Yang 2016), studies have used LLMs to generate detailed item descriptions (Acharya, Singh, and Onoe 2023), visual explanations (Tsai and Brusilovsky 2019), sentiment-oriented narratives (Chen, Yan, and Wang 2019), and interactive user-friendly justifications (Deldjoo 2024). Previous works have also shown the unique potential of natural language explanations in mitigating information overload and facilitating community integration (Chang, Harper, and Terveen 2016). This growing trend of work underscores the potential for LLM-driven recommendation strategy in strengthening the recommendation processes.
2.2 User Perceptions of Explanations
Users increasingly rely on recommender systems to navigate overwhelming volumes of information. Grounded in the Uses & Gratifications framework, individuals seek to satisfy their cognitive, affective, and utilitarian needs when engaging with social media content (Katz, Blumler, and Gurevitch 1973; Korhan and Ersoy 2016; Bashardoust, Feuerriegel, and Shrestha 2024). Accordingly, explanations that clarify why a recommendation is made can reduce cognitive effort, foster trust, and enhance decision confidence (Mandl et al. 2011; Jameson et al. 2015).
Extant research also identifies several factors shaping user perceptions of explanations. For instance, users may prefer human-generated recommendations, particularly for hedonic products (Wien and Peluso 2021). Explanations that reflect prior user feedback or behavior tend to resonate with individuals seeking tailored, value-reinforcing content (Lu et al. 2023). Moreover, individual differences in user profiles, such as cognitive engagement level, highlight the importance of detailed and transparent justifications for users who prioritize control and understanding (Chatti et al. 2022). Therefore, aligning explanation strategies with varied user motivations and perceptions is essential for designing effective and user-centric recommender systems.
2.3 Empirical User Studies on Explanations
Empirical research has analyzed how different explanations in recommender systems impact users (see Table 1). These studies typically use controlled experiments and surveys to evaluate diverse explanation strategies. Silva et al. (2024) conducted a user study (N=94) comparing ChatGPT-generated generic and personalized movie explanations. While both were well-received, personalization improved perceived effectiveness when recommendations were less relevant and enhanced persuasiveness in low-confidence scenarios. Similarly, Lubos et al. (2024) ran an online user study (N=97), finding that LLM-generated explanations outperformed template-based baselines in trust and satisfaction due to their contextual richness and fluency. In another user study of a fitness plan recommender system (N=341), Sun et al. (2023) showed that users trusted less of recommendations when explanations revealed reliance on their social-media friends’ activities, citing privacy and identity concerns. Instead, users preferred personalization based on their behavior. Together, these empirical findings provide insights into user perceptions and inform system design, but a comprehensive evaluation analyzing the cognitive, affective, and utilitarian needs is missing.
@@ 2.4 Research Gap
While the importance of recommendation explanations is widely acknowledged, research seldom examines how users interpret and respond to them. Furthermore, extant studies often rely on generic templates and have not fully explored the potential of LLM-generated contextualized explanations to address users’ cognitive, affective, and utilitarian needs (see Table 1). To address this gap, our study provides the first comprehensive user evaluation of LLM-generated contextualized explanations through the lens of the U&G framework. By examining how these explanations resonate with users’ multifaceted needs and influence their intentions, we aim to offer deeper insights into designing recommendation explanations that are more aligned with user motivations.
@@ 3.1 Overview
We conducted an online between-subjects experiment, where subjects (NN=759) were randomly assigned to one of three experimental conditions: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE).
@@ 3.2 Subjects
Initially, 1401 subjects completed the survey. We discarded responses from subjects failing an initial attention check. Subsequently, following prior research (Pennycook, Cannon, and Rand 2018) and our pre-registration, a subset of subjects were excluded for three reasons: (i) subjects who self-reported to have answered randomly, (ii) failed the second attention check, (iii) responded excessively fast (i.e., the duration was less than around 6 seconds for responding to the user perceptions). These users still received remuneration after the exclusion. Eventually, due to our strict exclusion criteria, 759 of the 1,401 subjects qualified, exceeding our estimated sample size of 750 to ensure the results achieved the desired statistical power.
@@ 3.3 Materials
We selected GPT-4 Turbo to generate movie recommendations for the main experiment. We conducted an additional experiment with Claude-3.5-Sonnet to validate the robustness of our study design and the generalizability of our findings (see Appendix C.1). We pre-compiled all recommendations and explanations in a corpus to ensure no variability in the outcomes across different LLM calls. The corpus is generated based on the pre-defined strategy and prompts below. Therefore, users with the same viewing history and assigned conditions received the identical recommendation explanation.
Recommendation Strategy
We implemented a sequential recommendation strategy following the algorithm in Table 5 in Appendix A. To reduce bias, users only rated movies they had not watched. For each user, we began the recommendation process with 10 distinct popular movies from the IMDb Top 250 list [^5] https://www.imdb.com/chart/top/. If the user had already watched a movie, we proceeded to new recommended movies until they reached an unwatched one. The recommended movies watched by users were also added to their viewing history and used to generate new recommendations. For example, if a user has watched Movie A before, LLMs will recommend Movie B. If the user has watched it as well, then LLMs will combine both movies in the explanation for Movie C. This process continued until either an unwatched recommendation was found or a maximum of 4 recommendations were generated.
Rather than relying on explicit subjective feedback (e.g., likes or dislikes), we collected information about whether users had previously watched each movie. This design simulates a cold-start scenario (Lam et al. 2008; Chaimalas et al. 2023), where prior viewing history is unavailable and where the system starts with general popular choices before adapting to observed behavior. This approach mirrors real-world conditions, where explicit user feedback is often sparse or unavailable.
Recommender websites (such as IMDb, Rotten Tomatoes, Google Maps, Yelp, and TripAdvisor) are typically user-independent and free of personalization, either because they have no access to user-specific information or personalization is unwanted. They only have access to consumption data from their registered users, that is, actual review contributors. For Yelp, it is estimated that only 0.67% of users are registered contributors, while the remaining users are visitors. [^6] See https://askwonder.com/research/percentage-yelp-users-mainly-write-reviews-restaurants-whrwmul7u, accessed March 12, 2025. Moreover, the main selling point of recommender websites like IMDb and Yelp is the comparison of products or services. As a result, users are not required to sign in to use the website. Forcing users to register on the website would probably be counterproductive and decrease activity, as users might not want to sign up due to privacy reasons. Consistent with these, we collect users’ viewing histories—rather than explicit preferences—to generate recommendations and explanations.
Prompt Generation
We followed previous practices (Giray 2023) when designing the prompts. Table 3 shows examples of contextualized and generic explanations. All explanations were limited to 100 words, aligning with the length of typical movie descriptions. To manage reading speed and cognitive load, each explanation was designed to include around three key features, such as the director, cast, plot, or scenes. Further, the length of the explanations was similar across the different conditions to ensure comparability.
@@
The main difference between CE (contextualized explanations) and GE (generic explanations) is in the explicit connection: while both explanations incorporate various elements (e.g., feature, popularity, similarity) to support the recommendations, the contextualized explanations naturally integrate the recommended movies with the user’s viewing history, explicitly highlighting how the user’s inputs shape the recommendation. Meanwhile, the generic explanations focus exclusively on the recommended movies. All generated movies were cross-checked with the IMDB dataset [^7] https://developer.imdb.com/non-commercial-datasets/ to ensure no hallucinations from the LLMs. Additionally, the corpus was validated by two experts with strong expertise in the field to ensure that all explanations used in the experiment were consistent and of fair quality.
@@ 3.4 Procedure
Afterward, we randomly reached out to NN=30 users for the follow-up interviews. [^8] Detailed interview questions in Appendix B. We integrated their responses to help interpret the quantitative analysis results and investigate how explanations affect user motivational needs.
@@ 3.5 Analyses
We conducted analyses based on user perceptions in the survey. Each participant could provide up to 10 ratings, resulting in a total of 2716 ratings from 759 participants in the main experiment (904 in NE, 911 in GE, and 901 in CE). We used hypothesis testing to identify differences in perceptions and consumption intentions across recommendation types, followed by mixed-linear regression to examine how user characteristics explain variation in both variables. Additionally, we conducted follow-up interviews to further interpret the quantitative results.
Hypothesis testing: We hypothesized that there are significant differences in the perceptions and consumption intentions between users receiving different recommendations. We conducted hypothesis testing on the four dependent variables of interest: Cognitive, Affective, Utilitarian, and UserIntent across the different conditions. Since the data did not meet the normality assumption (Lilliefors 1967), we used a Mann–Whitney test with Holm correction to identify pairwise differences between conditions (McHugh 2011). Later, we use $\mu$ to report the mean value.
@@
$Y_{ij}$ $=\beta_{0}+\beta_{1}\emph{Treatment}_{j}+\beta_{2}\emph{Age}_{j}+\beta_{3}\emph{Gender}_{j}+\beta_{4}\emph{Education}_{j}$
$\quad+\beta_{5}\emph{MovieConsumption}_{j}+\beta_{6}\emph{RecSysFrequency}_{j}$
$\quad+\beta_{7}\emph{RecSysUtility}_{j}+\beta_{8}\emph{HistoryNum}_{j}+\beta_{9}\emph{DisplayOrder}_{j}$
$\quad+\beta_{10}\emph{HistoryNum}_{j}\times\emph{MovieConsumption}_{j}$ (1)
$\quad+\beta_{11}\emph{HistoryNum}_{j}\times\emph{RecSysUtility}_{j}+u_{0j}+\epsilon_{ij}$
where ${Y}_{ij}$ denotes the dependent variables (i.e., Cognitive, Affective, Utilitarian, and UserIntent) for the $i$-th observation of the $j$-th subject, with intercepts $\beta_{0}$, subject-level random effects $u_{0j}$, coefficients $\beta_{1}$ through $\beta_{11}$, and an error term $\epsilon_{ij}$.